# Building 3D In-Context Learning Universal Model in Neuroimaging

### v1

@@ 4.2 Compared Models
Task-Specific Models. We compared Neuroverse3D against 3D task-specific models, which shared identical architecture and channel configurations but lacked the context branch. These task-specific models were trained on the held-out dataset, thereby mitigating domain shift concerns. This approach simulates the scenario where medical centers traditionally train models for specific scenario without employing ICL model. Identical data augmentation strategies were applied, excluding task augmentations that could potentially disadvantage task-specific models, such as the application of Sobel filters to segmentation masks. We evaluated performance under both few-shot learning scenarios, simulating performance trained with only context set data, and fully supervised learning scenarios, leveraging all available data from the meta-context set.
Other In-Context Learning Models. We implemented several other ICL methods for comparative analysis, including Painter [49], SegGPT [50], UniverSeg [11], and Neuralizer [13], all of which are designed for 2D image inputs. To accommodate the 3D context, we randomly sampled axial slices containing the segmentation target or brain to construct the 2D context set. The 2D context sizes were set to 1 for Painter [49], 8 for SegGPT [50], 32 for Neuralizer [13], and 64 for UniverSeg [11], corresponding to the reported optimal context sizes in their respective publications. The target 3D images were similarly decomposed into axial 2D slices and fed into these methods. The resulting 2D outputs were then reassembled into 3D images for metric evaluation. We downloaded the corresponding pretrained weights for these models to perform inference. UniverSeg [11] and SegGPT [50] are restricted to segmentation tasks, and all 2D models are incapable of performing 2D-to-3D task.
Comprehensive details regarding the training and evaluation protocols are provided in supplemental section C.
@@ 4.3 Model Comparison on Held-Out Datasets
In generation tasks, Neuroverse3D outperformed all other ICL models, surpassing Neuralizer and Painter across all tasks. Its performance closely matched few-shot models and nearly reached fully supervised levels in salt & pepper noise removal and skull stripping. This highlights the potential of 3D ICL models in generation tasks. Notably, our model’s performance gain in some generation tasks was less significant. This may be attributed to certain tasks, like modality transform and super-resolution, where 2D context sufficiently conveys task information, reducing the advantage of a 3D architecture.
Figure 5: Qualitative results for four representative tasks are presented. The tasks, from the first to the fourth row, are cerebral cortex segmentation, hippocampus segmentation, bias field correction, and Gaussian noise removal, respectively. To comprehensively illustrate the 3D image results, slices from both sagittal (first row) and axial (second row) views are displayed. Qualitative results for all tasks can be found in supplementary section C.3
Figure 6: Performance across different 3D context sizes. A group of Neuroverse3D-unseen models were trained to assess performance on tasks not encountered during training. Due to computational constraints, segmentation task evaluations here only include cerebral cortex, hippocampus, thalamus, and lateral ventricle. Generation task evaluations include bias field correction, Gaussian noise removal, and salt & pepper noise removal.
fig. 5 illustrates qualitative predictions of ICL models, showcasing results for both axial and sagittal slices. As 2D ICL models generate predictions on axial slices, discontinuities are frequently observed on sagittal slices. In segmentation, compared to Neuroverse3D, other models demonstrate lower accuracy on axial slices and frequent false positives/negatives in sagittal slices. In generation, Neuralizer generates excessive background errors, while Painter, despite reasonable performance on Gaussian denoising, struggles with neuroimage-specific tasks like bias correction. Conversely, Neuroverse3D maintains consistent accuracy across all tasks.
These results demonstrate Neuroverse3D’s accurate performance across diverse segmentation and generation tasks across different medical centers. This highlights the importance of 3D architectures for ICL models in neuroimaging, with Neuroverse3D offering a novel approach for building 3D universal models.
@@ 4.4 Impact of Context Size
Figure 6 illustrates model performance across varying context sizes. As expected, Neuroverse3D, along with other models, shows a significant performance increase with larger context sizes, emphasizing the critical role of context quantity for ICL models. Neuroverse3D exhibits the largest performance gain when the context size increases from 1 to 8. Furthermore, our model consistently outperforms other models across all context sizes in both segmentation and generation tasks. APSP overcomes memory limitations, enabling Neuroverse3D to support larger or even unlimited context sizes under limited resources, and further improve performance. SegGPT, due to memory limitations, could only support a maximum context size of 16.
@@ 4.5 Performance on Unseen Tasks
For segmentation tasks, Neuroverse3D-unseen exhibited a slight performance reduction compared to Neuroverse3D, with approximately a 5-point decrease in the Dice coefficient when the context size was greater than or equal to 8. However, it still significantly outperformed other ICL models. For generation tasks, Neuroverse3D-unseen showed a more substantial performance decline, although it remained marginally superior to Neuralizer. We attribute the greater performance drop in generation tasks compared to segmentation tasks in Neuroverse3D-unseen to the fact that, during training, Neuroverse3D-unseen encountered a wider variety of segmentation tasks, including segmentations of over thirty brain structures and random combinations of brain structures. This led to a learning process for segmentation tasks more close to meta-learning, enabling cross-task generalization. In contrast, the training set contained fewer generation tasks, leading the model to tend to memorize each generalization tasks. This suggests that incorporating a greater diversity of tasks into ICL models is a promising direction for further enhancing cross-task generalization capabilities.
@@ 4.6 Memory and Time Requirements
Table 7 presents a comparison of inference times across different models. Despite employing the most extensive context (8 $\times$ 128 2D slices), our model demonstrates superior inference speed compared to other models. This efficiency stems from the fact that 2D models process images slice-by-slice, necessitating the recomputation of context for each slice, thereby resulting in redundant context feature computations and prolonged inference durations. This further underscores the importance of developing 3D ICL models for neuroimaging.
@@ 4.7 Ablation Study of Loss Functions
We evaluated the model’s performance without the proposed modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss and gradient loss, as shown in table 1.
Removing the modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss led to a decline in segmentation performance, particularly in small, complex regions such as the hippocampus (see supplemental table 4 for details). This indicates that the modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss helps the model focus on more challenging regions, achieving a better balance in performance across various segmentation tasks.
@@
Table 1: Ablation study of loss functions demonstrating model performance across all segmentation and generation tasks, with results averaged over context sizes of 1, 2, 4, and 8.
@@ 5 Conclusion
This paper introduces Neuroverse3D, the first 3D ICL-based universal model for neuroimaging, addressing the critical challenge of high memory consumption. The proposed APSP and U-shaped fusion enable adaptive parallel-sequential context processing, supporting unlimited context. An optimized loss function balances performance across diverse tasks and enhances anatomical focus. Evaluated on cross-center data, Neuroverse3D significantly outperforms other ICL models in all tasks, matching the performance of fully supervised models in segmentation. Trained on large datasets, Neuroverse3D demonstrates robust generalization and eliminates retraining, showing strong practical potential. By overcoming memory limitations for ICL model, Neuroverse3D fundamentally paves the way for more diverse scenarios of ICL exploration such as processing of other organs in 3D.
@@ A.1 Gradient Equivalence
Here, we prove that the expected value of the gradient using Adaptive Parallel-Sequential Processing (APSP) is the same as the full-context gradient when no mini-contexts are discarded. The full-context gradient, exemplified by the $i$-th block output $\bar{c}_{\text{dec},j}^{i}$, with respect to parameters $\theta$ is defined as:
@@
where $L=\sum_{j=1}^{n}\text{len}(s_{j})$ ensures equal weighting of image-label pairs across mini-contexts. For brevity, we focus on the $i$-th block, with the understanding that the analysis applies analogously to other blocks.
@@
Equation (10) shows that the expected APSP gradient equals the full-context gradient $\nabla_{\theta}\mathcal{L}_{\text{full}}^{i}$. This analysis, exemplified for the $i$-th block, extends to all network blocks and parameters.
Thus, by employing the gradient of only the last mini-context and scaling by $n$, APSP preserves the full-context gradient expectation while enabling memory-efficient sequential computation.
@@ B Data
B.1 Data Sampling
@@
B.2 Image and Task Augmentation
@@ C.2 Evaluation
Performance was assessed using the Dice coefficient for segmentation tasks and the Peak Signal-to-Noise Ratio (PSNR) for generation tasks on held-out datasets. Each task and setting were evaluated 10 times with different context sets to compute the average and standard deviation, ensuring robust performance measurements.
@@ C.3 Qualitative Results
Figure 12: Qualitative results comparing of ICL models under a context size of 4.
Figure 13: Qualitative comparison with task-specific models under a context size of 4 and a few-shot training data size of 4.

### v2

@@ 4.2 Compared Models
Task-Specific Models. We compared Neuroverse3D against 3D task-specific models, which shared identical architecture and channel configurations but lacked the context branch. These task-specific models were trained on the held-out dataset, thereby mitigating domain shift concerns. This approach simulates the scenario where medical centers traditionally train models for specific scenario without employing ICL model. Identical data augmentation strategies were applied, excluding task augmentations that could potentially disadvantage task-specific models, such as the application of Sobel filters to segmentation masks. We evaluated performance under both few-shot learning scenarios, where models were trained with only context set data, and fully supervised learning scenarios, leveraging all available data from the meta-context set.
Other In-Context Learning Models. We compared our method with state-of-the-art ICL methods, including Painter [51], SegGPT [52], UniverSeg [12], and Neuralizer [14], all designed for 2D inputs. To accommodate the 3D context, we randomly sampled axial slices containing the segmentation target or brain to construct the 2D context set. The 2D context sizes were set to 1 for Painter [51], 8 for SegGPT [52], 32 for Neuralizer [14], and 64 for UniverSeg [12], corresponding to the reported optimal context sizes in their respective publications. The target 3D images were similarly decomposed into axial 2D slices and fed into these methods. The resulting 2D outputs were then reassembled into 3D images for metric evaluation. We downloaded the corresponding pretrained weights for these models to perform inference. UniverSeg [12] and SegGPT [52] are restricted to segmentation tasks, and all 2D models are incapable of performing 2D-to-3D reconstruction task.
Comprehensive details regarding the training and evaluation protocols are provided in supplemental Sec. C.
Figure 5: Qualitative results for four representative tasks are presented. To comprehensively illustrate the 3D image results, slices from both sagittal (first row) and axial (second row) views are displayed. Qualitative results for all tasks can be found in supplementary Sec. C.3
@@ 4.3 Model Comparison on Held-Out Datasets
In generation tasks, Neuroverse3D outperformed all other ICL models, surpassing Neuralizer and Painter across all tasks. Its performance closely matched few-shot models and nearly reached fully supervised levels in salt-and-pepper noise removal and skull stripping. This highlights the potential of 3D ICL models in generation tasks. Notably, our model’s performance gain in some generation tasks was less significant. This may be attributed to certain tasks, like modality transform and super-resolution, where 2D context sufficiently conveys task information, reducing the advantage of a 3D architecture.
Fig. 5 illustrates qualitative predictions of ICL models, showcasing results for both axial and sagittal slices. As 2D ICL models generate predictions on axial slices, discontinuities are frequently observed on sagittal slices. In segmentation, compared to Neuroverse3D, other models demonstrate lower accuracy and frequent false positives/negatives in sagittal slices. In generation tasks, Neuralizer generates excessive background errors, while Painter, despite reasonable performance on Gaussian denoising, struggles with neuroimage-specific tasks like bias correction and skull stripping. Conversely, Neuroverse3D maintains consistent accuracy across all tasks.
These results demonstrate Neuroverse3D’s accurate performance across diverse segmentation and generation tasks from different medical centers, attributed to its novel approach for 3D context processing and optimized loss function.
Figure 6: Performance across different 3D context sizes. A group of Neuroverse3D-unseen models was trained to assess performance on tasks not encountered during training. Due to computational constraints, segmentation task evaluations here only include cerebral cortex, hippocampus, thalamus, and lateral ventricle. Generation task evaluations include bias field correction, Gaussian noise removal, and salt-and-pepper noise removal.
@@ 4.4 Impact of Context Size
Figure 6 illustrates model performance across varying context sizes. Our model consistently outperforms other models across all context sizes in both segmentation and generation tasks. Besides, Neuroverse3D, along with other models, shows a significant performance increase with larger context sizes, emphasizing the critical role of context quantity for ICL models. The average metric improvement resulting from increased context size diminishes as the context size increases, reaching a plateau around 16. However, the standard deviation of predictions continues to decrease, leading to more stable predictions. APSP overcomes memory limitations, enabling Neuroverse3D to support larger or even unlimited context sizes under limited resources, thereby facilitating the practical use of Neuroverse3D. SegGPT, due to memory limitations, could only support a maximum context size of 16 in our experiments.
@@ 4.5 Performance on Unseen Tasks
For segmentation tasks, Neuroverse3D-unseen exhibited a slight performance reduction compared to Neuroverse3D, with approximately a 5-point decrease in the Dice coefficient after the context size was greater than or equal to 8. However, it still significantly outperformed other ICL models. For generation tasks, Neuroverse3D-unseen showed a more substantial performance decline, although it remained marginally superior to Neuralizer. We attribute the greater performance drop in generation tasks compared to segmentation tasks to the fact that, during training, Neuroverse3D-unseen encountered a wider variety of segmentation tasks, including segmentations of over thirty brain structures and random combinations of brain structures. This led to a learning process for segmentation tasks closer to meta-learning, enabling cross-task generalization. In contrast, the training set contained fewer generation tasks, leading the model to tend to memorize each generalization task. This suggests that incorporating a greater diversity of tasks into ICL models is a promising direction for further enhancing cross-task generalization capabilities.
@@ 4.6 Memory and Time Requirements
Supplemental Tab. 7 presents a comparison of inference times across different models. Despite employing the most extensive context (8 $\times$ 128 2D slices), our model demonstrates superior inference speed compared to other ICL models. This efficiency stems from the fact that 2D models process images slice-by-slice, necessitating the recomputation of context for each slice, thereby resulting in redundant context feature computations and prolonged inference durations. This further underscore the importance of developing Neuroverse3D which inherently avoid this issue.
@@ 4.7 Ablation Study of Loss Functions
We evaluated the model’s performance without the proposed modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss and gradient loss, as shown in Tab. 1.
Removing the modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss led to a decline in segmentation performance, particularly in small, complex regions such as the hippocampus (see supplemental Tab. 4 for details). This indicates that the modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss helps the model focus on more challenging regions, achieving a better balance in performance across various segmentation tasks.
@@
Table 1: Ablation study of loss functions demonstrating the model performance across all segmentation and generation tasks, with results averaged over context sizes of 1, 2, 4, and 8.
@@ 5 Conclusion
This paper introduces Neuroverse3D, the first 3D In-Context Learning (ICL) universal model for neuroimaging, addressing the critical challenge of high memory consumption inherent in ICL. The proposed APSP and U-shaped fusion enable adaptive parallel-sequential context processing, supporting unlimited context. An optimized loss function balances performance across diverse tasks and enhances anatomical focus. Evaluated on cross-center data, Neuroverse3D significantly outperforms other ICL models in all tasks, matching the performance of fully supervised models in segmentation. Trained on large datasets, Neuroverse3D demonstrates robust generalization and eliminates retraining, showing strong practical potential. By overcoming memory limitations for the ICL model, Neuroverse3D fundamentally paves the way for more diverse scenarios of ICL exploration such as processing other organs in 3D.
Acknowledgements
This work was supported in part by grants from the National Natural Science Foundation of P.R. China (62276081), Basic Foundation of Shenzhen Science and Technology National Key Program (CJGJZD20230724093959002), and The Major Key Project of PCL (PCL2023A09). Yanwu acknowledge the afffliation with the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
@@ A.1 Gradient Equivalence
Here, we prove that the expected value of the gradient using Adaptive Parallel-Sequential Processing (APSP) is the same as the full-context gradient when no mini-contexts are discarded. The full-context gradient, exemplified by the $i$-th stage output $\bar{c}_{\text{dec},j}^{i}$, with respect to parameters $\theta$ is defined as:
@@
where $L=\sum_{j=1}^{n}\text{len}(s_{j})$ ensures equal weighting of image-label pairs across mini-contexts. For brevity, we focus on the $i$-th stage, with the understanding that the analysis applies analogously to other stages.
@@
Equation (10) shows that the expected APSP gradient equals the full-context gradient $\nabla_{\theta}\mathcal{L}_{\text{full}}^{i}$. This analysis, exemplified for the $i$-th stage, extends to all network stages and parameters.
Thus, by employing the gradient of only the last mini-context and scaling by $n$, APSP preserves the full-context gradient expectation while enabling memory-efficient computation.
@@ B Data
B.1 Domain Shift Between Training and Test Data
We trained a discriminator to distinguish between images from the training set and those from the held-out set, achieving accuracy, precision, recall, and F1 scores of 0.934, 0.932, 0.918, and 0.925, respectively. Additionally, we trained a nnUNet model for brain anatomical segmentation using the training set. These results serve as proxy evidence for the presence of domain shift.
B.2 Data Sampling
@@
B.3 Image and Task Augmentation
@@ C.2 Evaluation
Performance was assessed using the Dice coefficient for segmentation tasks and the Peak Signal-to-Noise Ratio (PSNR) for generation tasks on held-out datasets. Each task and setting were evaluated 10 times with randomly selected context sets to compute the average and standard deviation, ensuring robust performance measurements.
@@ C.3 Qualitative Results
Figure 12: Qualitative results comparing of ICL models. In the SegGPT or Painter column, the results for segmentation tasks are from SegGPT, and the results for generation tasks are from Painter.
Figure 13: Qualitative comparison with task-specific models.
@@ C.4 Inference Cost and Model Size
C.5 Different Fusion Stage Configurations
Fusion Stage(s) 5 5–4 5–3 5–2 5–1 (Ours)
Segmentation (Dice) 0.8019 0.8118 0.8274 0.8347 0.8446
PET Segmentation (Dice) 0.2870 0.3162 0.3714 0.4945 0.5314
Generation (PSNR) 25.82 26.36 27.41 27.94 28.87
Table 8: Performance under different fusion stage configurations. Stage 5 corresponds to the deepest stage.
Tab. 8 summarizes the results of the ablation study on fusion stages. Denser connections improved the model’s performance, especially on PET segmentation—an unseen modality.
C.6 Impact of Different Context Modalities
Context Modality Bias Removal Gaussian Salt & Pepper Inpainting Super-res. 2D to 3D
T1 28.32 26.31 35.98 31.52 28.29 26.63
T2 24.93 18.88 34.86 29.18 25.34 14.15
CT 24.95 23.32 33.97 28.67 24.62 12.27
Table 9: Generation performance (PSNR) under different context modalities, with T1 as the target modality.
Tab. 9 presents the impact of different context modalities. For all tasks, performance is optimal when the context modality matches the target modality. The performance drop in salt-and-pepper noise removal is relatively small, potentially because this task is less dependent on contextual information. In contrast, Gaussian noise removal and 2D-to-3D transformation show more pronounced performance degradation, likely due to their greater reliance on context to convey task-relevant knowledge. These findings underscore the importance of modality alignment for the effective performance of ICL models.